OpenStudy (anonymous):

1/x + 1/sqrt(x) = 650 Find x. 1/x + 1/sqrt(x) = 650 Find x. @Mathematics

OpenStudy (zarkon):

1/625

OpenStudy (anonymous):

What are the steps?

OpenStudy (zarkon):

let u=1/sqrt(x) then u^2=1/x and your equation becomes \[u^2+u=650\] solve for u then for x